# Re-anchoring Quantum Monte Carlo with Tensor-Train Sketching

### v4

@@ 2. Preliminary
where each ${\phi}_{k}^{(n)}$ is a meanfield or separable state, $N$ is the number of walkers, and $e^{-\Delta\tau H}$ is decomposed as
@@
$e^{-\Delta\tau H}\widehat{\Psi}^{(n)}=\sum_{k}e^{-\Delta\tau H}{\phi}_{k}^{(n)}=\sum_{k}\sum_{\vec{x}}P(\vec{x})B(\vec{x}){\phi}_{k}^{(n)}\approx\sum_{k}{\phi}_{k}^{(n+1)}=:\widehat{\Psi}^{(n+1)},$
@@
$\phi^{(n+1)}_{k}=B(\vec{x}^{(n)}_{k}){\phi}_{k}^{(n)},\quad\vec{x}^{(n)}_{k}\sim P.$
As key algorithmic ingredients, cp-AFQMC further incorporates importance sampling when sampling $\vec{x}_{k}^{(n)}$in order to reduce the variance, which will also constrain the paths of the random walk when the sign problem is present. The starting point is a representation of the walker distribution of the following form
(5) $\widehat{\Psi}^{(n)}=\sum_{k=1}^{N}w_{k}^{(n)}\frac{\phi_{k}^{(n)}}{O_{\mathrm{tr}}(\phi^{(n)}_{k})},$
where,
$O_{\mathrm{tr}}(\phi_{k}^{(n)})=\max\left\{\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}},0\right\},$
and the weight $w_{k}^{(n)}$ is either positive, or
$w^{(n)}_{k}=0\ \text{if}\ O_{\mathrm{tr}}(\phi_{k}^{(n)})=0.$
With this representation, in cp-AFQMC it is possible to zero out some random walkers when $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}\leq 0$, hence constraining the paths.
To perform imaginary-time evolution, we incorporate a new probability distribution function $\widetilde{P}$, defined for a walker $\phi_{k}^{(n)}$ with positive weight $w_{k}^{(n)}>0$ as
(6) $\widetilde{P}_{k}^{(n)}(\vec{x})=\frac{1}{\mathcal{N}_{k}^{(n)}}\frac{O_{\mathrm{tr}}(B(\vec{x})\phi_{k}^{(n)})}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}P(\vec{x}),$
where $\mathcal{N}_{k}^{(n)}$ is the normalization constant. This new probability distribution function favors those walkers that have a large overlap with the trial wavefunction.
A remark here is that if an auxiliary field $\vec{x}$ leads to non-positive overlap with the trial function, so that $O_{\mathrm{tr}}(B(\vec{x})\phi_{k}^{(n)})=0$, then the probability is set to $\widetilde{P}_{k}^{(n)}(\vec{x})=0$, meaning that this $\vec{x}$ will not be sampled anymore. In this way, we only sample from those $\vec{x}$ such that the walker at the next step $B(\vec{x})\phi_{k}^{(n)}$ has a positive overlap with the trial function, and thus a positive weight. Therefore, the new walker $B(\vec{x})\phi_{k}^{(n)}$ is guaranteed to have a positive weight, as long as $\widetilde{P}_{k}^{(n)}(\vec{x})>0$ for at least one $\vec{x}$.
@@
(7) $e^{-\Delta\tau H}=\sum_{\vec{x}}P(\vec{x})B(\vec{x})\approx\sum_{\vec{x}:\widetilde{P}^{(n)}_{k}(\vec{x})>0}P(\vec{x})B(\vec{x})=\sum_{\vec{x}:\widetilde{P}^{(n)}_{k}(\vec{x})>0}\widetilde{P}^{(n)}_{k}(\vec{x})\left(\frac{P(\vec{x})}{\widetilde{P}^{(n)}_{k}(\vec{x})}B(\vec{x})\right).$
In a special case where $\widetilde{P}_{k}^{(n)}(\vec{x})>0$ for all $\vec{x}$, the approximation in (7) becomes exact. Plugging (7) into (5) gives
(8) $e^{-\Delta\tau H}\widehat{\Psi}^{(n)}$$\approx$ $\sum_{k}\sum_{\vec{x}:\widetilde{P}^{(n)}_{k}(\vec{x})>0}\widetilde{P}^{(n)}_{k}(\vec{x})w^{(n)}_{k}\frac{P(\vec{x})}{\widetilde{P}^{(n)}_{k}(\vec{x})}\frac{B(\vec{x}){\phi}_{k}^{(n)}}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}$
(9) $\approx$ $\sum_{k}w^{(n)}_{k}\frac{P(\vec{x}^{(n)}_{k})}{\widetilde{P}^{(n)}_{k}(\vec{x}^{(n)}_{k})}\frac{\phi_{k}^{(n+1)}}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}$
(10) $=$ $\sum_{k}\left(w^{(n)}_{k}\frac{P(\vec{x}^{(n)}_{k})}{\widetilde{P}^{(n)}_{k}(\vec{x}^{(n)}_{k})}\frac{O_{\mathrm{tr}}(\phi_{k}^{(n+1)})}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}\right)\frac{\phi_{k}^{(n+1)}}{O_{\mathrm{tr}}(\phi_{k}^{(n+1)})}$
(11) $=$ $\sum_{k}w_{k}^{(n+1)}\frac{\phi_{k}^{(n+1)}}{O_{\mathrm{tr}}(\phi_{k}^{(n+1)})},$
@@
$\phi_{k}^{(n+1)}=B(\vec{x}^{(n)}_{k}){\phi}_{k}^{(n)},\vec{x}^{(n)}_{k}\sim\widetilde{P}^{(n)}_{k},$
and
(12) $w^{(n+1)}_{k}=w^{(n)}_{k}\mathcal{N}_{k}^{(n)}.$
To estimate the ground-state energy, we use the mixed estimator
where $E_{\text{local}}$ represents the local energy of a walker:
@@
a Hamiltonian $H$, initial walkers $\widehat{\Psi}^{(0)}=\sum_{k}w_{k}^{(0)}\frac{\phi_{k}^{(0)}}{O_{\mathrm{tr}}(\phi_{k}^{(0)})}$, trial wavefunction $\Psi_{\mathrm{tr}}$, step size $\Delta\tau$, number of iteration steps $M$.
Walkers $\widehat{\Psi}^{(M)}=\sum_{k}w_{k}^{(M)}\frac{\phi_{k}^{(M)}}{O_{\mathrm{tr}}(\phi_{k}^{(M)})}$.
for $n=0,1,\cdots,M-1$ do
  • For walkers satisfying $\braket{\Psi_{\mathrm{tr}},\phi^{(n)}_{k}}>0$:
@@
      Sample $\vec{x}^{(n)}_{k}\sim\frac{1}{\mathcal{N}_{k}^{(n)}}\frac{O_{\mathrm{tr}}(B(\vec{x})\phi_{k}^{(n)})}{O_{\mathrm{tr}}(\phi_{k}^{(n)})}P(\vec{x})$, where $B(\vec{x}^{(n)}_{k})$ is defined by the decomposition $\exp(-\Delta\tau H)=\sum_{\vec{x}}P(\vec{x})B(\vec{x})$.
@@
      $\phi_{k}^{(n+1)}=B(\vec{x}^{(n)}_{k})\phi_{k}^{(n)}$.
    3. 3.
      $w_{k}^{(n+1)}=w_{k}^{(n)}\mathcal{N}_{k}^{(n)}$.
  • Periodically estimate the ground-state energy using the mixed estimator (13).
  • Periodically do population control as in Algorithm 5 in the appendix.
@@
In Appendix C.1, we detail the exact implementations for the transverse-field Ising (TFI) model.
@@ 2.2. Tensor-train sketching
where we denote $[n]=\{1,2,\cdots,n\}$ for some positive integer $n$. A TT representation is of the form
@@
We use the notation $m:n$ to denote the set $\{m,m+1,\cdots,n\}$ for two postive integers $m<n$, and further denote the truncated TT as
@@
which is a separable wavefunction. The advantage of TT is that it represents a $d$-dimensional tensor with a storage complexity linear in $d$. Moreover, the computational cost of the inner product of two TTs is also linear in $d$, which is numerically verified in Figure 12.
TT-sketching is a procedure to fit some given data into a low-rank TT. In this paper, the data type we consider is a sum of random walkers as in (5). Each walker is a separable state. A parallel method for this task is given [37], where a left sketching function $S=S_{1}\circ\cdots\circ S_{d}$ and a right sketching function $T=T_{1}\circ\cdots\circ T_{d}$ are used to compress the data. More specifically, denote
(14) $X_{j}(\xi_{j-1},\gamma_{j-1})=\sum_{i_{1:j-1}}\sum_{i_{j:d}}S_{1:j-1}(i_{1:j-1},\xi_{j-1})\widehat{\Psi}(i_{1:j-1},i_{j:d})T_{j:d}(i_{j},\gamma_{j-1}),$
and
$Y_{j}(\xi_{j-1},i_{j},\gamma_{j})=\sum_{i_{1:j-1}}\sum_{i_{j+1}:d}S_{1:j-1}(i_{1:j-1},\xi_{j-1})\widehat{\Psi}(i_{1:j-1},i_{j},i_{j+1:d})T_{j+1:d}(i_{j+1:d},\gamma_{j}).$
The tensor cores are then given by solving the equations
$\sum_{\gamma_{j-1}}X_{j}(\xi_{j-1},\gamma_{j-1})\widetilde{G}_{j}(\gamma_{j-1},i_{j},\gamma_{j})=Y_{j}(\xi_{j-1},i_{j},\gamma_{j}).$
This results in a high-rank TT with tensor cores $\widetilde{G}_{j}$. In practice, however, the matrices $X_{j}$ are usually low-rank. In this situation, we first apply a singular value thresholding to $X_{j}$, and then solve for the truncated tensor cores $G_{j}$, which amounts to trimming the TT to a lower rank $(r_{1},\ldots,r_{d-1})$.
For the sketching functions $S$ and $T$, we set them both to be TTs, with each tensor core $S_{j}$ and $T_{j}$ constructed as
(15) $S_{j}(\xi_{j-1},i_{j},\xi_{j})=\sum_{l}C_{S,j}(\xi_{j-1},l,\xi_{j})v_{l}(i_{j}),\ T_{j}(\xi_{j-1},i_{j},\xi_{j})=\sum_{l}C_{T,j}(\xi_{j-1},l,\xi_{j})v_{l}(i_{j}),$
@@
We aim to pick a relative small $\delta$ for the basis in (16), such that $\delta^{2}$ is still non-negligible but higher orders of $\delta$ can be ignored. Then $S$ and $T$ can be considered as $1$-cluster basis sketching functions [40, 41, 42].
In addition, instead of (14), we form the matrices $X_{j}$ sequentially by making use of the tensor cores we have already obtained, the detailed implementation of which is summarized in Algorithm 2. The tensor diagram illustrating the TT-sketching procedure is also provided in Figure 1. We remark that in this algorithm, as we sequentially update the tensor cores, the computation for forming $X_{j}$ can be reused, so the sketching algorithm has a linear time complexity in terms of the dimensionality $d$, which is verified in Figure 13 in the appendix.
Algorithm 2 TT-sketching
A wavefunction $\widehat{\Psi}(i_{1},\cdots,i_{d})$ in the form of (5); Left sketching function $S=S_{1}\circ\cdots\circ S_{d}$, right sketching function $T=T_{1}\circ\cdots\circ T_{d}$, with $S_{j}$ and $T_{j}$ constructed as in (15).
a rank-$r$ TT that approximates $\widehat{\Psi}$ with tensor cores being $G_{1},\cdots,G_{d}$.
First phase: construct an initial TT with tensor cores $\bm{\widetilde{G}_{1},\cdots,\widetilde{G}_{d}}$
  • For the first core $j=1$:
    1. 1.
      Form $Y_{1}$:
      $Y_{1}(i_{1},\gamma_{1})=\sum_{i_{2:d}}\widehat{\Psi}(i_{1};i_{2:d})T_{2:d}(i_{2:d},\gamma_{1});$
    2. 2.
      Solve for $\widetilde{G}_{1}$:
      $\widetilde{G}_{1}(i_{1},\gamma_{1})=Y_{1}(i_{1},\gamma_{1});$
  • For the cores $j=2,\cdots,d-1$:
    1. 1.
      Form $Y_{j}$:
      $Y_{j}(\xi_{j-1},i_{j},\gamma_{j})=\sum_{i_{1:j-1}}\sum_{i_{j+1:d}}S_{1:j-1}(i_{1:j-1},\xi_{j-1})\widehat{\Psi}(i_{1:j-1},i_{j},i_{j+1:d})T_{j+1:d}(i_{j+1:d},\gamma_{j});$
    2. 2.
      Form $X_{j}$:
      $X_{2}(\xi_{1},\gamma_{1})=\sum_{i_{1}}S_{1}(x_{1},\xi_{1})\widetilde{G}_{1}(i_{1},\gamma_{1}),\ j=2;$
      $X_{j}(\xi_{j-1},\gamma_{j-1})=\sum_{\xi_{j-2}}\sum_{\gamma_{j-2}}\sum_{i_{j-1}}X_{j-1}(\xi_{j-2},\gamma_{j-2})S_{j-1}(\xi_{j-2},i_{j-1},\xi_{j-1})\widetilde{G}_{j-1}(\gamma_{j-2},i_{j-1},\gamma_{j-1}),\ j>2;$
    3. 3.
      Solve for $\widetilde{G}_{j}$:
      $\sum_{\gamma_{j-1}}X_{j}(\xi_{j-1},\gamma_{j-1})\widetilde{G}_{j}(\gamma_{j-1},i_{j},\gamma_{j})=Y_{j}(\xi_{j-1},i_{j},\gamma_{j});$
  • For the last core $j=d$:
    1. 1.
      Form $Y_{d}$:
      $Y_{d}(\xi_{d-1},i_{d})=\sum_{i_{1:d-1}}S_{1:d-1}(i_{1:d-1},\xi_{d-1})\widehat{\Psi}(i_{1:d-1},i_{d});$
    2. 2.
      Form $X_{d}$:
      $X_{d}(\xi_{d-1},\gamma_{d-1})=\sum_{\xi_{d-2}}\sum_{\gamma_{d-2}}\sum_{i_{d-1}}X_{d-1}(\xi_{d-2},\gamma_{d-2})S_{d-1}(\xi_{d-2},i_{d-1},\xi_{d-1})\widetilde{G}_{d-1}(\gamma_{d-2},i_{d-1},\gamma_{d-1});$
    3. 3.
      Solve for $\widetilde{G}_{d}$:
      $\sum_{\gamma_{d-1}}X_{d}(\xi_{d-1},\gamma_{d-1})\widetilde{G}_{d}(\gamma_{d-1},i_{d})=Y_{d}(\xi_{d-1},i_{d});$
Second phase: compress the initial TT to targeted rank
  1. 1.
    Apply truncated SVD on $X_{j}$ by taking its $r_{j-1}$ largest singular values:
    $X_{j}(\xi_{j-1},\gamma_{j-1})\approx\sum_{\alpha_{j-1}=1}^{r_{j-1}}U_{j}(\xi_{j-1},\alpha_{j-1})\Sigma_{j}(\alpha_{j-1},\alpha_{j-1})V_{j}(\alpha_{j-1},\gamma_{j-1});$
  2. 2.
    Trim the tensor cores $\widetilde{G}_{1},\cdots,\widetilde{G}_{d}$ to $G_{1},\cdots,G_{d}$ with smaller size:
    $G_{j}(\alpha_{j-1},i_{j},\alpha_{j})=\sum_{\gamma_{j-1}}\sum_{\gamma_{j}}V_{j}(\alpha_{j-1},\gamma_{j-1})\widetilde{G}_{j}(\gamma_{j-1},i_{j},\gamma_{j})V_{j+1}(\alpha_{j},\gamma_{j}),2\leq j\leq d-1;$
    $G_{1}(i_{1},\alpha_{1})=\sum_{\gamma_{1}}\widetilde{G}_{1}(i_{1},\alpha_{1})V_{2}(\gamma_{1},\alpha_{1}),G_{d}(\alpha_{d-1},i_{d})=\sum_{\gamma_{d-1}}V_{d}(\alpha_{d-1},\gamma_{d-1})\widetilde{G}_{d}(\gamma_{d-1},i_{d});$
Figure 1. Tensor diagrams illustrating the TT-sketching described in Algorithm 2. As a demonstration, we present the steps involved constructing the third tensor core $G_{3}$ for a system with $d=6$. We refer the reader to [43] for a detailed introduction to tensor network diagrams.

### v5

@@ 2. Preliminary
where each ${\Phi}_{k}^{(n)}$, called a random walker, is a separable function, $N$ is the number of walkers. A discrete Hubbard-Stratonovich transformation [47] can be applied to decompose the operator $e^{-\Delta\tau H}$ as
@@
(5) $e^{-\Delta\tau H}\widehat{\Psi}^{(n)}=\sum_{k}e^{-\Delta\tau H}{\Phi}_{k}^{(n)}=\sum_{k}\sum_{\vec{x}}P(\vec{x})B(\vec{x}){\Phi}_{k}^{(n)}\approx\sum_{k}{\Phi}_{k}^{(n+1)}=:\widehat{\Psi}^{(n+1)},$
@@
$\Phi^{(n+1)}_{k}=B(\vec{x}^{(n)}_{k}){\Phi}_{k}^{(n)},\vec{x}^{(n)}_{k}\sim P.$
As key algorithmic ingredients, cp-AFQMC further incorporates importance sampling when sampling $\vec{x}_{k}^{(n)}$in order to reduce the variance. Given a walker $\Phi_{k}^{(n)}$ satisfying $\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}>0$, we incorporate a new probability distribution function $Q_{k}^{(n)}$, which is defined as
(6) $Q_{k}^{(n)}(\vec{x})=\frac{1}{\mathcal{N}_{k}^{(n)}}\frac{\max\{\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}},0\}}{\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}}P(\vec{x}),$
where $\mathcal{N}_{k}^{(n)}$ is the normalization factor. This new probability distribution function favors those walkers that have a large overlap with the trial wavefunction. Moreover, the walkers are guaranteed to have a positive overlap with the trial wavefunction, since the probability $Q_{k}^{(n)}(\vec{x})$ is set to 0 if some $\vec{x}$ leads to a non-positive overlap with $\Psi_{\mathrm{tr}}$, hence constraining the paths and causing a bias, which is also known as the sign problem. In an extreme case where all the auxiliary fields $\vec{x}$ lead to non-negative overlap $\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}\leq 0$, and thus $Q_{k}^{(n)}(\vec{x})=0$ for all $\vec{x}$, then this walker is eliminated from the procedure.
@@
(7) $e^{-\Delta\tau H}=\sum_{\vec{x}}P(\vec{x})B(\vec{x})\approx\sum_{\vec{x}:Q^{(n)}_{k}(\vec{x})>0}P(\vec{x})B(\vec{x})=\sum_{\vec{x}:Q^{(n)}_{k}(\vec{x})>0}Q^{(n)}_{k}(\vec{x})\left(\frac{P(\vec{x})}{Q^{(n)}_{k}(\vec{x})}B(\vec{x})\right).$
The approximate equality reflects the fact that the second sum runs over all the auxiliary fields $\vec{x}$, while the final expression restricts the summation to the subset of $\vec{x}$ satisfying $Q_{k}^{(n)}(\vec{x})>0$. In a special case where $Q_{k}^{(n)}(\vec{x})>0$ for all $\vec{x}$, the approximation in (7) becomes exact. With this new decomposition, the propagation (5) becomes
(8) $e^{-\Delta\tau H}\widehat{\Psi}^{(n)}\approx\sum_{k}\sum_{\vec{x}:Q^{(n)}_{k}(\vec{x})>0}Q^{(n)}_{k}(\vec{x})\left(\frac{P(\vec{x})}{Q^{(n)}_{k}(\vec{x})}B(\vec{x})\Phi_{k}^{(n)}\right)\approx\sum_{k}\Phi_{k}^{(n+1)}=:\widehat{\Psi}^{(n+1)},$
@@
(9) $\Phi_{k}^{(n+1)}=\frac{P(\vec{x}_{k}^{(n)})}{Q^{(n)}_{k}(\vec{x}_{k}^{(n)})}B(\vec{x}_{k}^{(n)})\Phi_{k}^{(n)},\vec{x}_{k}^{(n)}\sim Q_{k}^{(n)}.$
@@
a Hamiltonian $H$ and a corresponding decomposition $e^{-\Delta\tau H}=\sum_{\vec{x}}P(\vec{x})B(\vec{x})$, initial walkers $\widehat{\Psi}^{(0)}=\sum_{k}\Phi_{k}^{(0)}$, trial wavefunction $\Psi_{\mathrm{tr}}$, step size $\Delta\tau$, number of iteration steps $M$.
Walkers $\widehat{\Psi}^{(M)}=\sum_{k}\Phi_{k}^{(M)}$.
for $n=0,1,\cdots,M-1$ do For walkers satisfying $\braket{\Psi_{\mathrm{tr}},\Phi^{(n)}_{k}}>0$:
@@
    Sample an $\vec{x}^{(n)}_{k}$ from $Q_{k}^{(n)}$ defined in (6).
@@
    Update the walker $\Phi_{k}^{(n+1)}=\frac{P(\vec{x}_{k}^{(n)})}{Q^{(n)}_{k}(\vec{x}_{k}^{(n)})}B(\vec{x}_{k}^{(n)})\Phi_{k}^{(n)}$.
@@
We detail the exact implementations of Algorithm 1 and its application on the transverse-field Ising (TFI) model in the appendix.
@@ 2.2. Tensor-train sketching
where we denote $n]=\{1,2,\cdots,n\}$ for some positive integer $n$. A TT representation is of the form
@@
We use the notation $m:n$ to denote the set $\{m,m+1,\cdots,n\}$ for two positive integers $m<n$, and further denote the truncated TT as
@@
which is a separable wavefunction. The advantage of TT is that it represents a $d$-dimensional tensor with a storage complexity linear in $d$. Moreover, the computational cost of the inner product of two TTs is also linear in $d$. See numerical verification in the appendix.
A natural problem is how to fit given data into a low-rank TT. In this work, we consider data given as a sum of random walkers $\widehat{\Psi}$, as in (3). The wavefunction in (3) generally requires large storage, since the number of random walkers can be very large. Therefore, we aim to compress this representation into a low-rank TT. TT-sketching provides an effective procedure for accomplishing this task. The idea of TT-sketching is similar to randomized SVD [13] for matrix decomposition. More specifically, if the target tensor $\widehat{\Psi}$ has a low-rank TT structure, then the unfolding matrix $\widehat{\Psi}_{j}\in\mathbb{R}^{(n_{1}\cdots n_{j})\times(n_{j+1}\cdots n_{d})}$, defined by reshaping the tensor $\widehat{\Psi}$:
$\widehat{\Psi}_{j}(i_{1:j};i_{j+1:d})=\widehat{\Psi}(i_{1},\cdots,i_{d}),$
is also low-rank [27]. Thus, we expect that
$\mathrm{Range}(\widehat{\Psi}_{j})=\mathrm{Range}(\widehat{\Psi}_{j}T_{j}^{\mathrm{sketch}})$
with high probability, for some chosen sketching matrix $T_{j}^{\mathrm{sketch}}=[T_{\mathrm{sketch}}(i_{j+1:d};\xi_{j})]_{(i_{j+1:d};\xi_{j})}\in\mathbb{R}^{(n_{j+1}\cdots n_{d})\times R_{j}}$, where $R_{j}$ is the number of columns of the sketching matrix, usually selected to be larger than the target rank. Furthermore, we hope to construct tensor cores $G_{1},\cdots,G_{j}$ satisfying
$\mathrm{Range}(G_{1:j})=\mathrm{Range}(\widehat{\Psi}_{j}),$
where we treat $G_{1:j}=[G_{1:j}(i_{1:j};\xi_{j})]_{(i_{1:j};\xi_{j})}\in\mathbb{R}^{(n_{1}\cdots n_{j})\times R_{j}}$ as a matrix. To achieve these goals, we can simply require
$G_{1:j}=\widehat{\Psi}_{j}T_{j}^{\mathrm{sketch}}.$
The tensor diagram for these equations is shown in Figure 1, where we take a $4$-dimensional case as an illustration. We use $T=T_{1}\circ T_{2}\circ T_{3}\circ T_{4}$ in TT form as the sketching tensor.
Figure 1. Tensor diagrams illustrating the goal of TT-sketching using the 44-dimensional case as an example.
The equations in Figure 1 are over-determined and still have an exponentially large size. To further reduce computational cost, we introduce another sketching tensor $S=S_{1}\circ S_{2}\circ S_{3}\circ S_{4}$ on the left. The detailed implementation is illustrated in Figure 2, where each equation (except the first) can be written in the form
$\sum_{\xi_{j-1}}X_{j}(\gamma_{j-1},\xi_{j-1})G_{j}(\xi_{j-1},i_{j},\xi_{j})=Y_{j}(\gamma_{j-1},i_{j},\xi_{j}).$
The tensor cores can then be obtained sequentially by solving these equations via multiplying the pseudoinverse of $X_{j}$ with $Y_{j}$. After obtaining each tensor core $G_{j}$, we can further trim it to the target size by inserting a projector $V_{j}V_{j}^{T}$, where $V_{j}\in\mathbb{R}^{R_{j}\times r_{j}}$ and $r_{j}$ is the target rank. We group $G_{j}V_{j}$ as the final tensor core, and multiply $V_{j}^{T}$ into the next tensor core.
Figure 2. Tensor diagrams illustrating the detailed implementation of TT-sketching using the 44-dimensional case as an example.
For the sketching tensors $S$ and $T$, we construct their tensor cores $S_{j}$ and $T_{j}$ as
$S_{j}(\xi_{j-1},i_{j},\xi_{j})=\sum_{l}C_{S,j}(\xi_{j-1},l,\xi_{j})v_{l}(i_{j}),\ T_{j}(\xi_{j-1},i_{j},\xi_{j})=\sum_{l}C_{T,j}(\xi_{j-1},l,\xi_{j})v_{l}(i_{j}),$
@@
We aim to pick a relative small $\delta$ for the basis in (10), such that $\delta^{2}$ is still non-negligible but higher orders of $\delta$ can be ignored. Then $S$ and TT can be considered as $1$-cluster basis sketching functions [50, 2, 28].
We remark that in this algorithm, as we sequentially compute the tensor cores, the left-hand-side tensors in Figure 2 can also be computed sequentially, meaning that the computational costs for obtaining each tensor core are the same. Thus, the sketching algorithm has a linear time complexity in terms of the dimensionality $d$. See numerical verification in the appendix.